OpenStudy (anonymous):

For the reaction, NH3(g) + HCl(g) NH4Cl(s) run at 25°C, ΔG = −91368 J and ΔS = −284 J/K. What is the enthalpy change for this reaction? 184,632 J −91,084 J −98,468 J −176,000 J A negative change in free energy, or −ΔG, represents a spontaneous reaction. ^ this is true

OpenStudy (anonymous):

25°C = 298 K Use ΔG = ΔH - T ΔS ΔH= ΔG + TΔS = −91368 + (298)(−284)=−91368-84632=-175730 Therefore ΔH is approx - 176,000 J

OpenStudy (anonymous):

A negative change in free energy, or −ΔG, represents a spontaneous reaction. ^ this is true

OpenStudy (anonymous):

For a reaction to be spontaneous, ΔG should be negative. So, now decide yourself :)

OpenStudy (anonymous):

oh okay so i am right.

OpenStudy (anonymous):

But If a reaction experiences a large decrease in entropy and absorbs a large amount of heat, the reaction will be spontaneous?

OpenStudy (anonymous):

most likely it will be non-spontaneous. However it would be unsafe to comment on this. I would rather use this formula ΔG = ΔH - T ΔS and see if ΔG is positive or negative if ΔG is positive --> non-spontaneous reaction if ΔG is negative --> spontaneous reaction

OpenStudy (anonymous):

okay last one The free energy of an object is the total available energy that the object possesses. this is one is false? correct?

OpenStudy (anonymous):

Well, can you give me a a better word for free energy

OpenStudy (anonymous):

In other words I don't understand the term "free energy" in this context.